OpenStudy (monica0597):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. How do you solve this? I have a few problems like this and i want to know how to solve them instead of asking answers for each one.

OpenStudy (anonymous):

You can use the equation again. y - y1 = m(x - x1)

OpenStudy (monica0597):

thanks

OpenStudy (monica0597):

y-y1=m(x-x1) y-(-3)=6x-(-2) What do i do here now? Did i even get the steps right?

OpenStudy (anonymous):

Solve

OpenStudy (monica0597):

how did you get the 6 to get 36?

hero (hero):

@monica0597, you should post the list of answer choices first. The reason you should do this is because the correct answer can be expressed in several different forms.

OpenStudy (monica0597):

y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

hero (hero):

Okay, so those answer choices are in slope-intercept form. Are you familiar with the general formula for slope intercept?

OpenStudy (monica0597):

yeah, y=mx+b. i started solving it but i know i made a mistake somewhere. i don't know where though...

hero (hero):

Okay, now think about this. You are given one of the points on the line and the slope: (x,y) = (-2,-3) m = -6 So basically, you have three values you can insert into the general formula: x = -2 y = -3 m = -6 If you do this, the only thing missing is b. If you insert x, y, and m into the formula and solve for b, you'll be nearly finished with the problem. Try inserting the given values into the formula. Let me know how it looks afterwards.

OpenStudy (monica0597):

after i inserted the values i got this:  y-(-3)=6(x-(-2))